A speed trap is made by placing two distant radar cameras, 150 m apart, along the highway. Suppose you are driving, and you notice the speed trap and begin slowing down the instant you cross the first radar camera. If you are moving at a rate of 42 m/s and the speed limit is 35 m/s. What must your minimum acceleration be for you to have an average speed within the speed limit by the time your car crosses the second radar camera?
1
Expert's answer
2021-10-12T12:31:58-0400

"t=s/v_0=150/35=4.28\ (s)"


"s=vt-at^2/2\to a=2(s-vt)/t^2=2(42\cdot4.28-150)/4.28^2\approx3.25\ (m/s^2)" . Answer